### v2

@@ Abstract
This article presents the automatic checking of research outputs package ACRO, which assists researchers and data governance teams by automatically applying best-practice principles-based statistical disclosure control (SDC) techniques on-the-fly as researchers conduct their analyses. ACRO distinguishes between: research output that is safe to publish; output that requires further analysis; and output that cannot be published because it creates substantial risk of disclosing private data. This is achieved through the use of a lightweight Python wrapper that sits over well-known analysis tools that produce outputs such as tables, plots, and statistical models. This adds functionality to (i) identify potentially disclosive outputs against a range of commonly used disclosure tests; (ii) apply disclosure mitigation strategies where required; (iii) report reasons for applying SDC; and (iv) produce simple summary documents trusted research environment staff can use to streamline their workflow. The major analytical programming languages used by researchers are supported: Python, R, and Stata. The ACRO code and documentation are available under an MIT license at https://github.com/AI-SDC/ACRO
@@
Data privacy, data protection, data science, privacy, Python, R, Stata, statistical disclosure control.
@@ 1 INTRODUCTION
Statistical agencies and other custodians of secure facilities such as trusted research environments (TREs) [1] provide researchers with access to confidential data under the ‘Five Safes’ data governance framework [2]. This enforces five orthogonal layers of safety procedures with the last requiring explicit checking of research outputs for disclosure risk. This can be a time-consuming and costly task, requiring skilled staff. This article discusses the development of a free and open source tool for automating the statistical disclosure control (SDC) of routine research outputs such as tables, plots, and statistical models. The goal is to make the clearance process more efficient and timely, and to allow the skilled checkers to focus their attention on the less straightforward cases.
More specifically, ACRO (automatic checking of research outputs) assists researchers and output checkers by distinguishing between research output that is safe to publish, output that requires further analysis, and output that cannot be published because of substantial disclosure risk. This is achieved through the use of Python’s capacity to override standard commands for creating tables, regressions, and other queries. This keeps the syntax identical while simultaneously augmenting analysis commands by running and reporting appropriate disclosure risk assessments. A schematic illustration of the ACRO workflow is shown in Figure LABEL:fig:schematic.
The operational design philosophy is extensively documented by Green et al. [3] who studied the characteristics that an automated SDC solution needs to have to be feasible, effective, and a positive choice for users. The essential criteria are that it should be:
@@
    Free and open source: ACRO is available on GitHub under an MIT License at: https://github.com/AI-SDC/ACRO.
@@
    Easy to obtain: in addition to the source code available on GitHub, a Python package is available through the Python Package Index (PyPI), which provides a simple management system for distributing Python libraries and facilitating continuous updates. Similarly, the R ACRO package [4] is available through the Comprehensive R Archive Network (CRAN). ACRO is cross-platform, supporting GNU/Linux, macOS, and Windows; with v0.4.6 supporting Python 3.8–3.12.
@@
    Easy to use: ACRO reduces barriers to adoption via a front-end application programming interface (API) that is identical to those already commonly used by researchers in their favoured language. Researchers prepare their data and statistical queries in the usual way, in their preferred language. When they are satisfied with the format of a query that generates, for example, a table or a regression, they then repeat it using the same commands prefixed by acro. The lightweight translation functions then call the Python back-end, which executes the queries and performs the requisite output checks. Easy to understand help and documentation is available within the GitHub repository.
@@
    Comprehensive: ACRO identifies potentially disclosive outputs with a range of commonly used disclosure tests, applies mitigation strategies where required, and flags outputs for human review that cannot be definitively classified as safe or unsafe. The current version of ACRO uses suppression and rounding to protect privacy, which are the most common forms of disclosure mitigation strategy.
@@
    Transparent: applying disclosure mitigation where required, but with clear and rapid explanations that enable (i) researchers to understand and revise their requests; and (ii) an auditable record for TRE staff with simple summary reports to streamline their workflow; ACRO provides researchers with immediate feedback by displaying the results of both the checks and the output to the researcher, with the option for automatic suppression where appropriate. Details of the queries and results are stored in a list, which may subsequently be written to file for review by a human output checker. ACRO gives researchers control over the outputs that are submitted for review, e.g., the removal of unwanted outputs and choice of output format; currently JSON [5] or Microsoft Excel®.
@@
    Able to support exceptions under principles-based regimes; ACRO enables researchers to add comments to requested outputs and ensures that researchers supply supporting evidence for all outputs that cannot be classified as safe.
@@
    Able to implement an organisation’s business rules for primary and secondary disclosure, which may vary across datasets or users. Within ACRO, SDC rules and parameters can be uniquely configured through a simple YAML [6] file specifying the risk appetite.
@@
ACRO is a product of the “Semi-Automated Checking of Research Outputs” (SACRO) project funded by Data and Analytics Research Environments UK (DARE UK). [^1] https://dareuk.org.uk/driver-project-sacro/ This work builds upon a previous Eurostat-funded project, in which Green et al. [7] developed a proof-of-concept prototype for the proprietary Stata software. As part of the SACRO project, the SACRO team committed to review and re-develop the theory and operational guidelines for output SDC. The aim was threefold; first, to bring together key points from the SDC literature (and fill in some of the theoretical gaps) to provide an integrated guide to both theory and practice of output checking; second, to develop a new approach to output SDC based on classifications into groups; for details, see Derrick et al. [8]; third, to explicitly link theory to operational rules and their implementation in manual and automatic checking regimes.
Five UK TREs (OpenSAFELY [^2] https://www.opensafely.org at the University of Oxford and four Scottish Safe Havens) were involved in the SACRO project to provide detailed feedback on user and output checker perspectives; OpenSAFELY also took the lead in the design of a graphical user interface (GUI). This group also directly tested the feasibility of installing and allowing the Python code to run on their systems as TREs differ in their perceptions of Python’s ‘riskiness’. Moreover, the SACRO team contacted a large number of TREs in the UK and abroad, and set up a network of interested parties willing to be testers. Several engagement events with this group identified how they worked and what they would expect from an automatic solution.
Subsequently, a ‘Community of Interest’ group has been established to provide on-going peer support and new feature prioritisation. Currently, the mailing list333mailto:sdc-reboot@jiscmail.ac.uk has 50 subscribers from 25 different organisations. To illustrate the breadth of uptake, just some of those who are deploying SACRO and encouraging its use by researchers include: eDRIS/Public Health Scotland and the Scottish National Safe Haven; Clinical Practice Research DataLink (Medicines & Healthcare Products Regulatory Agency, England), and the Secure Anonymised Information Linkage (SAIL) Databank (Wales).
The remainder of this article is organised as follows. Section 2 describes the current state-of-the-art for output SDC with the Five Safes framework and SDC implementations. Section 3 details the SDC checks implemented with ACRO, as well as how to install and use the Python, R, and Stata packages. Section 4 describes the GUI for viewing and managing the results of SDC produced by ACRO. Finally, Section 5 discusses future plans.
@@ 2.1 FIVE SAFES AND OUTPUT DISCLOSURE CONTROL
The Five Safes framework [9] is a set of principles that enable services to provide safe research access to their data and has been adopted by a range of TREs, including the UK Office for National Statistics (ONS), Health Data Research UK (HDR UK), and the National Institute for Health Research Design Service (NIHR), as well as many others worldwide.
@@
Mitigation strategies commonly suppress, round, or add noise to ‘unsafe’ outputs in order to prevent disclosure [14]. For example, a minimum threshold rule applied to the number of observations used by a statistic ensures that there is sufficient uncertainty with respect to any individual respondent. Dominance rules protect large respondent values from being approximated where the contribution to a statistic is dominated by only a few individuals. For example, the $p$%-rule sorts the $N$ observations by magnitude and checks whether the sum of the smallest $N-3$ observations is at least $p$of the largest observation. The $NK$ rule checks that the largest $N$ observations contribute less than $K$of the total. Not all aggregation statistics are considered safe; typically, reporting minima or maxima values of a subgroup are prohibited.
Regressions are generally protected by checking that the residual degrees-of-freedom exceeds a minimum threshold. The purpose is to prevent an equation from masquerading as a model. Residual degrees of freedom is defined as the number of observations minus the number of restrictions embodied in the test or model. For example, for a simple linear regression it is $N-K$, where $K$ is the number of coefficients including the intercept; for chi-square it is $N$ test degrees of freedom. Following Brandt et al. [15], 10 is typically used as the minimum.
Devising rules for each individual statistic used by analysts is infeasible, however it is possible to combine statistics into groups based not on statistical relation, but on common disclosure risks and solutions. Consequently, Ritchie et al. [16] propose the use of ‘statistical barns’ (statbarns) as a means to classify statistics for disclosure control purposes. A statbarn is a collection of statistics that share the same characteristics for disclosure control purposes. That is, their mathematical form is similar; they share the same risks; they share the same responses to those risks; and output checking rules are applicable to all. For example, pie charts, histograms, and scatter plots are all forms of frequency table. Table LABEL:table:statbarns lists the 14 statbarns that have been identified, with 12 classified for output checkers.
@@ 2.3 PRINCIPLES-BASED DISCLOSURE CONTROL
Traditional approaches to SDC within TREs have focused on rules-based approaches, where a rule is a hard limit with no exceptions permitted. More recently, it has been recognised that a principles-based approach to SDC is more effective since defining hard rules that are neither too restrictive nor too loose is extremely difficult [17]. For example, a frequency table with zeros in some cells is highly disclosive if they reveal that no one in a given region earned more (or less) than $k$ USD. In contrast, reporting that no one born without organ $k$ was diagnosed with cancer of said organ would usually be recognised as a “structural zero”, and permitted.
@@
Semi-automated output checking therefore aims to improve the rigour and consistency of the output disclosure control process and reduce human workload by automatically applying best-practice principles-based SDC techniques. That is, automatically identifying, reporting, and optionally applying mitigation strategies to disclosive outputs where possible. Outputs are categorised as ‘safe’ or ‘unsafe’. ‘Safe’ outputs requiring no or minimal further changes can be expedited through the clearing process, whereas ‘unsafe’ outputs can be prioritised for human review [13].
@@
  • Improved auditing by generating outputs in a structured format.
@@
  • Increased consistency of decision making within and between TREs.
@@ 2.4 TOOLS FOR SDC
The Research Data and Service Centre (RDSC) of the Deutsche Bundesbank provide two tools to allow researchers to check whether the results generated in their research project comply with the RDSC’s SDC rules [21]. Researchers run a special command directly after generating their results and get immediate feedback. The RDSC provide nobsdes5/nobsreg5, which are a set of Stata ado files, and the sdcLog [22] package, which provides a set of new R functions. These tools help researchers in checking descriptive statistics and models, and in calculating extreme values that are not individual data; however these packages have not been updated since March 2022. Since these tools implement new commands, researchers must learn them and adapt their workflow.
With the aim of developing a more general semi-automated solution for use by researchers in TREs, improving the efficiency of the process, and (where applicable) reducing the amount of user training required, a recent Eurostat project [7] developed a proof-of-concept prototype in Stata where primary disclosure is regulated by a set of simple rules. Here we build upon the lessons learned from developing that prototype with the aim of expanding its coverage and impact whilst retaining a minimal learning curve. We achieve this by adopting the principles-based approach to SDC outlined by Ritchie et al. [16] and develop the ACRO toolkit, which is primarily implemented in Python, but with cross-language support for R and Stata, using analytical interfaces familiar to researchers.
3 THE ACRO TOOLKIT

### v3

@@ Abstract
This article presents a free and open source toolkit that supports the semi-automated checking of research outputs (SACRO) for privacy disclosure within secure data environments. SACRO is a framework that applies best-practice principles-based statistical disclosure control (SDC) techniques on-the-fly as researchers conduct their analyses. SACRO is designed to assist human checkers rather than seeking to replace them as with current automated rules-based approaches. The toolkit is composed of a lightweight Python package that sits over well-known analysis tools that produce outputs such as tables, plots, and statistical models. This package adds functionality to (i) automatically identify potentially disclosive outputs against a range of commonly used disclosure tests; (ii) apply optional disclosure mitigation strategies as requested; (iii) report reasons for applying SDC; and (iv) produce simple summary documents trusted research environment staff can use to streamline their workflow and maintain auditable records. This creates an explicit change in the dynamics so that SDC is something done with researchers rather than to them, and enables more efficient communication with checkers. A graphical user interface supports human checkers by displaying the requested output and results of the checks in an immediately accessible format, highlighting identified issues, potential mitigation options, and tracking decisions made. The major analytical programming languages used by researchers (Python, R, and Stata) are supported by providing front-end packages that interface with the core Python back-end. Source code, packages, and documentation are available under MIT license at https://github.com/AI-SDC/ACRO
@@
Data privacy, data protection, privacy, statistical disclosure control, statistical software.
@@ 1 INTRODUCTION
Statistical agencies and other custodians of secure data facilities such as trusted research environments (TREs) [1] provide researchers with access to confidential data under the ‘Five Safes’ data governance framework [2]. This enforces five orthogonal layers of safety procedures with the last requiring the explicit checking of research outputs for disclosure risk. This can be a time-consuming and costly task, requiring skilled staff, and is currently impeding efforts to scale-up the use of TREs [3].
Existing solutions have attempted to automate output statistical disclosure control (SDC) with rules-based approaches. However, defining hard rules that should apply regardless of context is extremely difficult [4] and fully-automated solutions are unable to address a number of potential vulnerabilities such as secondary disclosure through differencing attacks.
This article discusses the development of a free and open source toolkit for semi-automating the SDC of routine research outputs such as tables, plots, and statistical models (SACRO). SACRO implements a principles-based SDC methodology, which aims to assist rather than replace human checkers, and ensure that researchers have met minimum required standards before submitting their request for release. The goal is to make the clearance process more efficient and timely, and to allow the skilled checkers to focus their attention on the less straightforward cases.
The SACRO toolkit is composed of (i) a Python package (ACRO) that implements a number of automated checks and optional mitigation strategies that researchers use on-the-fly as they conduct their analyses; and (ii) a graphical user interface (GUI) that uses the reports generated with ACRO to assist checkers in reviewing outputs, tracking decisions, and providing an auditable record (SACRO viewer). Additional languages such as R and Stata are supported with wrapper packages that interface with the Python ACRO back-end.
More specifically, the ACRO tool assists researchers and output checkers by distinguishing between research output that requires further analysis, and output that cannot be published because of a substantial risk of disclosing private data. The tool produces summaries that indicate whether the automated tests passed, failed, or require review. A status of “pass” does not indicate that the tool has certified that an output is safe for publication, but that it has passed a minimum level of checks that staff normally would have had to manually perform and provides information to help expedite further review.
The details of the checks and any mitigations are summarised for skilled checkers to assist in their review. This is achieved through the use of Python’s capacity to override standard commands for creating tables, regressions, and other queries. This keeps the syntax identical while simultaneously augmenting analysis commands by running and reporting appropriate disclosure risk assessments. A schematic illustration of the SACRO workflow is shown in Figure LABEL:fig:schematic.
A significant advantage of the semi-automated principles-based approach is that researchers maintain control over the SDC process and can select the most appropriate mitigation strategies with approval from checkers, as opposed to the strict and brittle automatic rules-based approaches that are currently in use [4].
The operational design philosophy is extensively documented by Green et al. [5] who studied the characteristics that a semi-automated SDC solution needs to have to be feasible, effective, and a positive choice for users. The essential criteria are that it should be:
@@
    Free and open source: the SACRO tools are available on GitHub under MIT License at: https://github.com/AI-SDC/ACRO.
@@
    Easy to obtain: in addition to the source code available on GitHub, a Python ACRO package is available through the Python Package Index (PyPI), which provides a simple management system for distributing Python libraries and facilitating continuous updates. Similarly, the R ACRO package [6] is available through the Comprehensive R Archive Network (CRAN). ACRO and the SACRO viewer GUI are cross-platform, supporting GNU/Linux, macOS, and Windows; with v0.4.7 supporting Python 3.9–3.13.
@@
    Easy to use: SACRO reduces barriers to adoption via a front-end application programming interface (API) that is identical to those already commonly used by researchers in their favoured language. Researchers prepare their data and statistical queries in the usual way, in their preferred language. When they are satisfied with the format of a query that generates, for example, a table or a regression, they then repeat it using the same commands prefixed by acro. The lightweight translation functions then call the Python back-end, which executes the queries and performs the requisite output checks. Easy to understand help and documentation is available within the GitHub repository.
@@
    Comprehensive: ACRO identifies potentially disclosive outputs with a range of commonly used disclosure tests, applies mitigation strategies as requested, and flags outputs that have failed any tests. The current version of ACRO implements suppression and rounding as options for researchers to protect privacy, which are the most common forms of disclosure mitigation strategy.
@@
    Transparent: identifying and optionally applying disclosure mitigation where requested, but with clear and rapid explanations that enable (i) researchers to understand and revise their requests; and (ii) an auditable record for TRE staff with simple summary reports to streamline their workflow; ACRO provides researchers with immediate feedback by displaying the results of both the checks and the output to the researcher, with the option to apply mitigations where appropriate. Details of the queries and results are stored in a list, which may subsequently be written to file for review by a human output checker. ACRO gives researchers control over the outputs that are submitted for review, e.g., the removal of unwanted outputs and choice of output format; currently JSON [7] or Microsoft Excel®.
@@
    Able to support exceptions under principles-based regimes; ACRO enables researchers to add comments to requested outputs and ensures that researchers supply supporting evidence for all outputs that fail any of the tests.
@@
    Able to implement an organisation’s business rules for primary and secondary disclosure, which may vary across datasets or users. Within ACRO, SDC heuristics and parameters can be uniquely configured through a simple YAML [8] file specifying the risk appetite.
@@
The SACRO toolkit is a product of the “Semi-Automated Checking of Research Outputs” project funded by Data and Analytics Research Environments UK (DARE UK). [^1] https://dareuk.org.uk/driver-project-sacro/ This work builds upon a previous Eurostat-funded project, in which Green et al. [9] developed a proof-of-concept prototype for the proprietary Stata software. As part of the SACRO project, the SACRO team committed to review and re-develop the theory and operational guidelines for output SDC. The aim was threefold; first, to bring together key points from the SDC literature (and fill in some of the theoretical gaps) to provide an integrated guide to both theory and practice of output checking; second, to develop a new approach to output SDC based on classifications into groups; for details, see Derrick et al. [10]; third, to explicitly link theory to operational rules and their implementation in manual and automatic checking regimes.
Five UK TREs (OpenSAFELY [^2] https://www.opensafely.org at the University of Oxford and four Scottish Safe Havens) were involved in the SACRO project to provide detailed feedback on user and output checker perspectives; OpenSAFELY also took the lead in the design of the GUI. This group also directly tested the feasibility of installing and allowing the Python code to run on their systems as TREs differ in their perceptions of Python’s ‘riskiness’. Moreover, the SACRO team contacted a large number of TREs in the UK and abroad, and set up a network of interested parties willing to be testers. Several engagement events with this group identified how they worked and what they would expect from a semi-automatic solution.
Subsequently, a ‘Community of Interest’ group has been established to provide on-going peer support and new feature prioritisation. Currently, the mailing list333mailto:sdc-reboot@jiscmail.ac.uk has over 50 subscribers from more than 25 different organisations. To illustrate the breadth of uptake, just some of those who are deploying SACRO and encouraging its use by researchers include: eDRIS/Public Health Scotland and the Scottish National Safe Haven; Clinical Practice Research DataLink (Medicines & Healthcare Products Regulatory Agency, England), and the Secure Anonymised Information Linkage (SAIL) Databank (Wales).
The remainder of this article is organised as follows. Section 2 discusses the current state-of-the-art for output SDC with the Five Safes framework and SDC implementations. Section 3 describes the SACRO toolkit, including the SDC checks implemented, as well as how to install and use the packages. Section 4 provides details of the GUI for viewing and managing the results of SDC produced by ACRO. Finally, Section 5 discusses future plans.
@@ 2.1 FIVE SAFES AND OUTPUT DISCLOSURE CONTROL
The Five Safes data governance framework [11] is a set of principles that enable services to provide safe research access to their data and has been adopted by a range of TREs, including the UK Office for National Statistics (ONS), Health Data Research UK (HDR UK), and the National Institute for Health Research Design Service (NIHR), as well as many others worldwide.
@@
Five mitigation strategies are frequently employed to protect ‘unsafe’ outputs: (i) generalisation, which reduces the specificity of a sensitive value; (ii) suppression, which hides sensitive values completely; (iii) permutation, where records are partitioned and the values of sensitive attributes are shuffled within subgroups; (iv) perturbation, where sensitive values are replaced with synthetically generated data; and (v) anatomisation, where sensitive values are split into separate tables to break any associations [15].
Commonly, a minimum threshold rule is applied to the number of observations used by a statistic to ensure that there is sufficient uncertainty with respect to any individual respondent [16].
Dominance rules protect large respondent values from being approximated where the contribution to a statistic is dominated by only a few individuals. For example, the $p$%-rule sorts the $N$ observations by magnitude and checks whether the sum of the smallest $N-3$ observations is at least $p$of the largest observation. The $NK$ rule checks that the largest $N$ observations contribute less than $K$of the total. Not all aggregation statistics are considered safe; typically, reporting minima or maxima values of a subgroup are prohibited.
Regressions are generally protected by checking that the residual degrees-of-freedom exceeds a minimum threshold. The purpose is to prevent an equation from masquerading as a model. Residual degrees of freedom is defined as the number of observations minus the number of restrictions embodied in the test or model. For example, for a simple linear regression it is $N-K$, where $K$ is the number of coefficients including the intercept; for chi-square it is $N$ test degrees of freedom. Following Brandt et al. [17], 10 is typically used as the minimum. Categorical data is also frequently protected by enforcing minimum counts of the sample sizes, and applying rounding rules based on the significant digits.
Devising rules for each individual statistic used by analysts is infeasible, however it is possible to combine statistics into groups based not on statistical relation, but on common disclosure risks and solutions. Consequently, Ritchie et al. [18] propose the use of ‘statistical barns’ (statbarns) as a means to classify statistics for disclosure control purposes.
A statbarn is a collection of statistics that share the same characteristics for disclosure control purposes. That is, their mathematical form is similar; they share the same risks; they share the same responses to those risks; and output checking rules are applicable to all. For example, pie charts, histograms, and scatter plots are all forms of frequency table. Table LABEL:table:statbarns lists the 14 statbarns that have been identified, with 12 classified for output checkers.
@@ 2.3 PRINCIPLES-BASED DISCLOSURE CONTROL
Traditionally, TREs have focused on rules-based approaches to output SDC, where a rule is a hard limit with no exceptions permitted. More recently, it has been recognised that a principles-based approach to SDC is more effective since defining hard rules that are neither too restrictive nor too loose is extremely difficult [4]. For example, a frequency table with zeros in some cells is highly disclosive if they reveal that no one in a given region earned more (or less) than $k$ USD. In contrast, reporting that no one born without organ $k$ was diagnosed with cancer of said organ would usually be recognised as a “structural zero”, and permitted.
@@
Semi-automated output checking therefore aims to improve the rigour and consistency of the output disclosure control process and reduce human workload by automatically applying best-practice principles-based SDC techniques. That is, automatically identifying, reporting, and optionally applying mitigation strategies to disclosive outputs where possible.
@@
  • Improved auditing by generating structured outputs.
@@
  • Increased consistency within and between TREs.
@@ 2.4 TOOLS FOR SDC
Additionally, Statistics Netherlands have developed sdcMicro [21], which is an R package and GUI that provides various methods for anonymising data and performing risk estimation.
@@
The Research Data and Service Centre (RDSC) of the Deutsche Bundesbank provide two tools to allow researchers to check whether the results generated in their research project comply with the RDSC’s SDC rules [23]. Researchers run a special command directly after generating their results and get immediate feedback. The RDSC provide nobsdes5/nobsreg5, which are a set of Stata ado files, and the sdcLog [24] package, which provides a set of new R functions. These tools help researchers in checking descriptive statistics and models, and in calculating extreme values that are not individual data. However, since these tools implement new commands, researchers must learn them and adapt their workflow.
With the aim of developing a more general semi-automated solution for use by researchers in TREs, improving the efficiency of the process, and (where applicable) reducing the amount of user training required, a recent Eurostat project [9] developed a proof-of-concept prototype in Stata where primary disclosure is regulated by a set of simple rules. Importantly, the design philosophy was of like-for-like replacement using the same call signatures as common Stata commands, but prefixed by a special keyword. Here we build upon the lessons learned from developing that prototype with the aim of expanding its coverage and impact whilst retaining a minimal learning curve. We achieve this by adopting the principles-based approach to SDC outlined by Ritchie et al. [18] and develop the SACRO toolkit, which is primarily implemented in Python, but with cross-language support for R and Stata, using analytical interfaces familiar to researchers.
3 THE SACRO TOOLKIT
3.1 OVERVIEW
The SACRO toolkit consists of (i) the ACRO Python package, which implements a range of automated disclosure tests and optional mitigation algorithms; (ii) additional language packages that create their own Python virtual environment and provide wrapper functions that interface with the ACRO Python package; and (iii) a GUI that uses the reports generated to assist staff checkers in performing their reviews. Since the additional language packages use the same Python back-end for running and reporting the tests, the results are consistent across different platforms and languages. In addition, the same configuration files can be used across different languages and the reports generated for use with the GUI are identical.
As the additional language packages depend on the core Python back-end, they are pinned to a specific release of the Python ACRO package. This enables updates to be made to the Python package, including potentially breaking API changes, without affecting dependant languages. The additional language packages are then updated to incorporate these changes.
Since the SACRO toolkit has been designed as a drop-in replacement in order to maintain familiar interfaces, it necessarily depends on particular versions of the analysis tools supported. These versions are similarly pinned to specific ranges of versions to maintain stability and interoperability, and kept under continuous review. The versions currently supported are noted within the documentation.
A CHANGELOG markdown file is included in the Python GitHub repository which maintains summaries of all updates made for each version of the tool; and a NEWS markdown file similarly records changes made for the R package.